### v1

# HumanoidBench: Simulated Humanoid Benchmark for Whole-Body Locomotion and Manipulation

Abstract
Humanoid robots hold great promise in assisting humans in diverse environments and tasks, due to their flexibility and adaptability leveraging human-like morphology. However, research in humanoid robots is often bottlenecked by the costly and fragile hardware setups. To accelerate algorithmic research in humanoid robots, we present a high-dimensional, simulated robot learning benchmark, HumanoidBench, featuring a humanoid robot equipped with dexterous hands and a variety of challenging whole-body manipulation and locomotion tasks. Our findings reveal that state-of-the-art reinforcement learning algorithms struggle with most tasks, whereas a hierarchical learning baseline achieves superior performance when supported by robust low-level policies, such as walking or reaching. With HumanoidBench, we provide the robotics community with a platform to identify the challenges arising when solving diverse tasks with humanoid robots, facilitating prompt verification of algorithms and ideas. The open-source code is available at https://sferrazza.cc/humanoidbench_site.
Benchmark Dexterous hands Action dim. DoF Task horizon # Tasks Skills11footnotemark: 1
[…]
1PnP: Pick-and-place / P: Push / I: Insert / R: Reach / Po: Pose / IR: In-hand re-orientation / H: Hold / L: Lift / Ro: Rotate / Lo: Locomotion / BM: Whole-body (humanoid) Manipulation / St: Stabilization
TABLE I: Comparison of simulated robotic manipulation benchmarks. Our humanoid robot benchmark tests a variety of complex, long-horizon task with a large action space.
I Introduction
Humanoid robots have long held promise to be seamlessly deployed in our daily lives. Despite the rapid progress in humanoid robots’ hardware (e.g., Boston Dynamics Atlas, Tesla Optimus, Unitree H1), their controllers are fully or partially hand-designed for specific tasks, which requires significant engineering efforts for each new task and environment, and often demonstrates only limited whole-body control capabilities.
In recent years, robot learning has shown steady progress in both robotic manipulation [10, 61, 12] and locomotion [23, 63]. However, scaling learning algorithms to humanoid robots is still challenging and has been delayed mainly due to such robots’ costly and unsafe real-world experimental setups.
To accelerate the progress of research for humanoid robots, we present the first-of-its-kind humanoid robot benchmark, HumanoidBench, with a diverse set of locomotion and manipulation tasks. Our simulated humanoid benchmark demonstrates a variety of challenges in addressing learning for autonomous humanoid robots, such as the intricate control of robots with complex dynamics, sophisticated coordination among various body parts, and addressing long-horizon complex tasks, while providing an accessible, fast, safe, and inexpensive testbed to robot learning researchers.
HumanoidBench contributes (1) a simulation environment comprising a humanoid robot with two dexterous hands, as illustrated in LABEL:fig:teaser; (2) a variety of tasks, spanning locomotion, manipulation, and whole-body control, incorporating humans’ everyday tasks; (3) a standardized benchmark to evaluate the progress of the community on high-dimensional humanoid robot learning and control; and finally, (4) benchmarking results of the state-of-the-art reinforcement learning (RL) algorithms and a hierarchical RL approach.
The simulation environment of HumanoidBench uses the MuJoCo [52] physics engine. For the simulated humanoid robot, we mainly opt for a Unitree H1 humanoid robot [^1] https://www.unitree.com/h1, which is relatively affordable and offers accurate simulation models [58], with two dexterous Shadow Hands [^2] https://www.shadowrobot.com/dexterous-hand-series/ attached to its arms. Our environment can easily incorporate any humanoid robots and end effectors; thus, we provide other models, including Agility Robotics Digit [^3] https://agilityrobotics.com/robots and the Robotiq 2F-85 gripper.
The HumanoidBench task suite includes $15$ distinct whole-body manipulation tasks involving a variety of interactions, e.g., unloading packages from a truck, wiping windows using a tool, catching and shooting a basketball. In addition, we provide $12$ locomotion tasks (not requiring hands’ dexterity), which can serve as primitive skills for whole-body manipulation tasks and provide a set of easier tasks to verify algorithms. The benchmarking results on this task suite show how the state-of-the-art RL algorithms struggle with controlling the complex humanoid robot dynamics and solving the most challenging tasks, illustrating ample opportunities for future research.
II Related Work
Deep reinforcement learning (RL) has made rapid progress with the advent of standardized, simulated benchmarks, such as Atari [4] and continuous control [6, 50] benchmarks. In robotic manipulation, most existing simulated environments are limited to quasi-static, short-horizon skills, having focused on tasks like picking and placing [6, 21, 62, 56, 32], in-hand manipulation [42, 38, 7], and screwing [37].
Complex manipulation tasks, such as block stacking [11], kitchen tasks [14], and table-top manipulation [22, 34, 29], have been introduced but are still limited to a combination of pushing, picking, and placing. On the other hand, the IKEA furniture assembly environment [27], BEHAVIOR [48, 28], and Habitat [49] present diverse long-horizon (mobile) manipulation tasks, with their main focus being on high-level planning by abstracting complex low-level control problems, while FurnitureBench [19] introduces a simulated benchmark for complex long-horizon furniture assembly tasks with sophisticated low-level control. However, most of these benchmarks use a single-arm manipulation setup with either a parallel gripper or a dexterous hand [8, 42], limiting the types of object interactions and not addressing the challenges of coordinating multiple parts of a body [26], e.g., multiple fingers, arms, and legs.
Robosuite [62] includes a handful of bimanual manipulation tasks, while more recently [9] and [59] have introduced additional benchmarks that require coordinating two floating robot hands, i.e., not attached to any arm base. While bimanual manipulation is one of the key objectives of humanoid robots, most benchmarks in humanoid research have so far focused on the locomotion challenges of such platforms [7, 24, 39, 3]. In this regard, such simulations have accelerated research on control algorithms [5, 40, 41, 35], ultimately leading to achieve robust humanoid locomotion in the real world [1, 43].
Recent works have extended humanoid simulations to different domains involving a certain degree of manipulation, i.e., tennis [60], soccer [16], ball manipulation [53] and catching [33], and box moving [55]. However, all these works focus on demonstrating their approaches on specific humanoid tasks and lack a diversity of tasks. In addition, most of the previous work focuses on simplistic humanoid models [33, 53], leading to inaccurate physics and collision handling. This motivates us to implement a comprehensive simulated humanoid benchmark based on real-world hardware and consisting of a diverse set of whole-body control tasks with careful design choices for diversity and usability.
Figure 0: Example egocentric visual (top-left) and whole-body tactile (right) observations when the humanoid interacts with a package in the truck environment. In the right figure, the two cameras on the robot head are highlighted in green, while continuous tactile pressure readings are indicated in shades of red and yellow. Note that here, for ease of visualization, we are not showing shear forces and tactile readings on the back of the robot, which are also implemented in our environment.
(a) push
(b) cabinets
(c) high_bar
(d) door
(e) truck
(f) cubes
(g) bookshelf
(h) basketball
(i) window
(j) spoon
(k) kitchen
(l) package
(m) powerlift
(n) room
(o) insert
Figure 1: HumanoidBench manipulation task suite. We devise $15$ benchmarking whole-body manipulation tasks that cover a wide variety of interactions and difficulties. This figure illustrates an initial state (left) and target state (right) for each task.
(a) walk
(b) stand
(c) run
(d) reach
(e) hurdle
(f) crawl
(g) maze
(h) sit
(i) balance
(j) stairs
(k) slides
(l) pole
Figure 2: HumanoidBench locomotion task suite. We devise $12$ benchmarking locomotion tasks that cover a wide variety of interactions and difficulties. This figure illustrates an initial state (left) and target state (right) for each task, except for walk.
III Simulated Humanoid Robot Environment
In this section, we describe our simulated environment and discuss relevant design choices for the simulated humanoid robot. As illustrated in Figure , we use the Unitree H1 humanoid robotFootnote 1 with two dexterous Shadow HandsFootnote 2 as the primary robotic agent of our benchmark. We simulate this humanoid robot using MuJoCo [52] adapting the Unitree H1 model provided by Unitree [^4] https://github.com/unitreerobotics/unitree_ros and the dexterous Shadow Hand models available through MuJoCo Menagerie [58].
Humanoid Body. We implement Unitree H1Footnote 1 and Agility Robotics DigitFootnote 3, which are well-known humanoid robots with their model files freely available [58]. Unitree H1 is primarily used in our benchmark as we observed faster learning, which we ascribe to a simpler mechanical design compared to Agility Robotics Digit, which features passive joints actuated through a four-bar linkage.
Dexterous Hands. We use two dexterous Shadow HandsFootnote 2, which also have model files freely available [58], and have shown impressive manipulation capabilities both in simulation [59] and the real world [2]. To make the simulated robot have more human-like morphology, we remove the cumbersome forearms of the dexterous Shadow Hands in HumanoidBench. While this is not currently a realistic model, we anticipate the trend in the industry towards developing slimmer, human-like hands (e.g. Tesla Optimus, Figure 01) so that our design choice aligns better with next-generation humanoid robots. In addition, we also provide models for the Robotiq 2F-85 parallel-jaw gripper and the hands available in the Unitree collectionFootnote 4.
Observations. Our simulated environment supports the following observations:
  • Proprioceptive robot state (i.e. joint angles and velocities) and task-relevant environment observations (i.e. object poses and velocities).
  • Egocentric visual observations from two cameras placed on the robot head (see Figure ).
  • Whole-body tactile sensing using the MuJoCo tactile grid sensor (see Figure ). We design tactile sensing at the hands with high resolution and in other body parts with low resolution, similar to humans, with a total of $448$ taxels spread over the entire body, each providing three-dimensional contact force readings. Similar distributed force readings have been captured on real-world systems both on humanoid bodies [36] and end-effectors [46]. The implementation of such spatially distributed contact sensing required non-trivial mesh adaptations and refinements, which we detail in the appendix.
Although other sensory inputs are available from the environment, to investigate challenges in whole-body control of humanoid robots, we first focus on the state-based environment setup, where proprioceptive robot states and object states are used for the agent’s input in HumanoidBench. We leave extending our environment to benchmarking multimodal perception capabilities [57, 47] of humanoid robots as future work. Note that in our state-based environment, we essentially maintain the robot observations the same across tasks to minimize domain knowledge, in contrast to tailoring it to the specific tasks [51].
[…]
(a) walk
(b) stand
(c) run
(d) reach
(e) hurdle
(f) crawl
(g) maze
(h) sit_simple
(i) sit_hard
(j) balance_simple
(k) balance_hard
(l) stair
(m) slide
(n) pole
Figure 3: Learning curves of RL algorithms (locomotion). The curves are averaged over three random seeds and the shaded regions represent the standard deviation. The dashed lines qualitatively indicate task success. We run PPO on the walk task but it is not visible in the plot since it only achieves very low returns.
(a) push
(b) cabinet
(c) highbar
(d) door
(e) truck
(f) cube
(g) bookshelf_simple
(h) bookshelf_hard
(i) basketball
(j) window
(k) spoon
(l) kitchen
(m) package
(n) powerlift
(o) room
(p) insert_small
(q) insert_normal
Figure 4: Learning curves of RL algorithms (manipulation). The curves are averaged over three random seeds and the shaded regions represent the standard deviation. The dashed lines qualitatively indicate task success. Note that kitchen is the only environment with a purely discrete, sparse reward, with a maximum of $4$.
IV HumanoidBench
Humanoid robots promise to solve human-like tasks in human-tailored environments, possibly using human tools. However, their form factor and hardware challenges make real-world research challenging, making simulation a crucial tool to advance algorithmic research in the field.
To this end, we present HumanoidBench, a humanoid benchmark for robot learning and control, which features a high-dimensional action space (up to $61$ different actuator) and enables research in complex whole-body coordination.
We benchmark $27$ tasks, consisting of $12$ locomotion tasks and $15$ distinct manipulation tasks, as illustrated in Figure 2 and Figure 1. A set of locomotion tasks aim to provide interesting but simpler humanoid control scenarios, bypassing intricate dexterous hand control. On the other hand, whole-body manipulation tasks render a comprehensive evaluation of the state-of-the-art algorithms on challenging tasks with unique challenges that require coordination across the entire robot body, ranging from toy examples (e.g. pushing a box on a table) to practical applications (e.g. truck unloading, shelf rearrangement).
Below we briefly describe the tasks that are part of the benchmark. Further details about each of the tasks, including task initialization and reward functions, are provided in the appendix.
IV-A Locomotion Tasks
[…]
  • stand: Maintain a standing pose throughout the provided amount of time.
[…]
  • reach: Reach a randomly initialized 3D point with the left hand.
[…]
  • maze: Reach the goal position in a maze by taking multiple turns at the intersections.
  • sit: Sit onto a chair situated closely behind the robot.
  • balance: Stay balanced on the unstable board.
[…]
  • pole: Travel in forward direction over a dense forest of high thin poles, without colliding with them.
IV-B Whole-Body Manipulation Tasks
  • push: Move a box to a randomly initialized 3D point on a table.
  • cabinet: Open four different types of cabinet doors (e.g. hinge doors, sliding door, drawer).
  • highbar: Athletically swing while staying attached to a horizontal high bar until reaching a vertical upside-down position.
  • door: Pull a door and traverse it while keeping the door open.
  • truck: Unload packages from a truck by moving them onto a platform.
  • cube: Manipulate two cubes in-hand until they both reach a randomly initialized target orientation.
  • bookshelf: Pick and place several items across shelves in a given order.
  • basketball: Catch a ball coming from random directions and throw it into the basket.
  • window: Grab a window wiping tool and keep its tip parallel to a window by following a prescribed vertical velocity.
  • spoon: Grab a spoon and use it to follow a circular pattern inside a pot.
  • kitchen [14]: Execute a sequence of actions in a kitchen environment, namely, open a microwave door, move a kettle, and turning burner and light switches.
  • package: Move a box to a randomly initialized target position.
  • powerlift: Lifting a barbell shaped object of a designated mass.
[…]
  • insert: Insert the ends of a rectangular peg into two tight target blocks.
V Benchmarking Results
To identify the challenges in learning with humanoid robots, we benchmark reinforcement learning (RL) algorithms on HumanoidBench, which promises for robots to learn from their own experience. Remarkably, this class of algorithms requires limited domain expertise and does not necessarily rely on expert demonstrations, which are not only expensive but also challenging to collect for humanoid robots.
V-A Baselines
We evaluate all tasks in our benchmark with four reinforcement learning methods (DreamerV3, TD-MPC2, SAC, PPO). Please refer to appendix, Appendix C, for implementation details.
  • DreamerV3 [17]: the state-of-the-art model-based RL algorithm, learning from imaginary model rollouts.
  • TD-MPC2 [18]: the state-of-the-art model-based RL algorithm with online planning.
  • SAC (Soft Actor-Critic [15]): the state-of-the-art off-policy model-free RL algorithm.
  • PPO (Proximal Policy Optimization [45]): the state-of-the-art on-policy model-free RL algorithm.
V-B Results
We report benchmarking results in Figure 3 and Figure 4, where we ran each of the algorithms for approximately 48 hours, resulting in the visible differences in environment steps (e.g. $2$M steps for TD-MPC2, $10$M steps for DreamerV3). We only run PPO on a subset of tasks (walk, kitchen, door, package), given its inferior performance without massive parallelization. Each of the environments is evaluated with a combination of dense rewards and sparse subtask completion rewards, and for each of these we provide qualitative measures of task success (see dashed lines in Figure 3 and Figure 4).
All the baseline algorithms perform below the success threshold on most tasks, particularly struggling on tasks that require long-horizon planning and intricate whole-body coordination in a high-dimensional action space. Surprisingly, these state-of-the-art RL algorithms require a large number of steps to learn even simple locomotion tasks, such as walk, which has been extensively studied with a simplified humanoid agent in the DeepMind Control Suite [51].
This poor performance is mainly attributed to the high-dimensionality of the state and action spaces of our humanoid robot agent with dexterous hands. Although the hands of the humanoid robot are barely used for most locomotion tasks, the RL algorithms fail to ignore these information, which makes policy learning challenging. In addition, these high-dimensional state and action spaces result in a much larger exploration space, which makes exploration slow or infeasible with simple maximum entropy approaches. This implies the need for incorporating behavioral priors or common sense knowledge about the world that can ease the exploration problem, when it comes to learning more complex agents, like humanoid robots. We investigate this problem further in Section V-C.
This problem becomes even more severe in manipulation tasks, resulting in particularly low rewards in all manipulation tasks. Before learning any manipulation skills, an agent must learn locomotion skills to balance and move towards an object or the world to interact. All the policies barely learn to stabilize using the dense reward, but struggle to learn any complex manipulation skills.
(a) walk
(b) push
Figure 5: Performance with and without dexterous hands. The curves are averaged over three random seeds and the shaded regions represent the standard deviation.
V-C With Hands vs. Without Hands
As discussed in Section V-B, controlling humanoid robots with dexterous hands is challenging due to their high degrees of freedom and complex dynamics. Thus, we investigate the difficulty of RL training with a large action space (i.e. additional $42$-DoFs with two dexterous Shadow Hands) on walk that does not necessarily require to control dexterous hands. The results in Figure 5 show that the presence of hands, with their additional joints and actuators, leads to a large decrease in performance compared to training the same task without the dexterous hands.
We perform an additional experiment to further confirm whether such difficulties stem from the dimensionality of the action space. In particular, we benchmark our full robot model, but fix the actuation of the hands, which we set to zero. In this way, the action dimensionality is reduced from $61$ in the original model to $19$, but the additional observations and masses induced by the presence of the hands are retained. On this experiment, the algorithms only slightly decrease their performance (see Blocked in the figure), confirming that most of the performance drop is indeed due to the increased action dimensionality. We repeat the same experiment, with similar results, by also removing the hand joints from the observation space (see Reduced in the figure). We observe similar trends in the more complex push task, which however, also presents substantially different dynamics in the task approach (e.g. pushing with and without hands).
V-D Flat vs. Hierarchical Reinforcement Learning
Figure 6: Our hierarchical architecture, where a robust low-level reaching policy (trained using PPO) follows targets provided by a high-level planning policy.
As shown above, flat, end-to-end RL approaches fail to learn most of the tasks in HumanoidBench. The results above show that even state-of-the-art RL algorithms struggle on the most complex tasks in the benchmark. Many of such tasks require long-horizon planning and necessitate acquiring a diverse set of skills (e.g., balancing, walking, reaching, etc.) to successfully achieve the desired objective. We argue that these issues can be mitigated by introducing additional structure into the learning problem. In particular, we explore a hierarchical learning paradigm, where one or multiple low-level skill policies are provided to a high-level planning policy that sends setpoints to lower-level policies, see Figure 6. In practice, such setpoints comprise the action space of the high-level policy. This framework is very general, and there are no constraints on how to obtain both low-level and high-level policies. However, here we focus on training both of these through reinforcement learning.
We report results of such a hierarchical approach on two different tasks, namely, the push and package tasks. As a low-level skill, push uses a one-hand reaching policy, which allows the robot to reach a 3D point in space with its left hand, while package uses a two-hand reaching policy, where both hands are commanded to reach different 3D targets.
Ideally, the low-level skill policies are pretrained blocks that can be reused across tasks. In particular, our hierarchical architecture considers the low-level policy as a frozen block that does not improve during training of the high-level policy. As such, the low-level reaching policies in our push and package tasks need to be very robust to cope with the continually shifting reaching targets that the high-level policy sets during exploration. However, the results in the previous section show how even a one-hand reaching task is hard to learn for the algorithms benchmarked in this work.
On a separate note, while our experiments above confirm that the on-policy PPO exhibits poor sample efficiency compared to the other off-policy algorithms, it is worth noting that PPO has achieved significant success in robotic locomotion by exploiting large-scale parallelization of environments on GPUs [31]. We exploit hardware acceleration using the recently released MuJoCo MJX [^5] https://mujoco.readthedocs.io/en/stable/mjx.html, which enables training PPO on thousands of parallel environments. We employ a simplified H1 model that only considers collisions between feet and ground in our MJX experiments, as in our experience the advantages stemming from parallelization are largely reduced when considering all numerous humanoid geometries (hindering training of the more complex benchmark tasks via MJX). We also remove the hands from the model to further increase training efficiency. We modify the reaching task by resetting the target during each episode whenever this has been reached. We also apply force perturbations at each of the links during training. We train the one-hand reaching policy for $2$ billion steps ($36$ hours) and the two-hand reaching policy for 4 billion steps ($60$ hours) on a total of $32,768$ parallel environments. As a result of the robustness introduced by the force perturbations, our model directly transfers to the full humanoid model (with original collision meshes) on the classical MuJoCo simulation, and is also robust to the additional masses introduced by the presence of dexterous hands. We provide the MuJoCo MJX training code as part of our software package.
(a) push
(b) package
Figure 7: Comparison between flat policies and hierarchical policies. The curves are averaged over three random seeds and the shaded regions represent the standard deviation.
As mentioned above, we use the reaching policies obtained through MJX training as low-level policies of our hierarchical architecture. For the push and package tasks, we restrict the range of reaching targets to the robot workspace to facilitate exploration, and the high-level policies are trained with either DreamerV3 or TDMPC2. Results are shown in Figure 7, where the hierarchical architecture significantly outperforms the end-to-end baselines on the push task, achieving very high success rates. While the low-level policy has undergone additional pretraining, this can be in principle reused across tasks. We note a less pronounced performance improvement in the more challenging package task. While getting closer to picking up the package, the policy still struggles in lifting it (having never experienced it during training), which limits the overall task performance.
These results confirm that the tasks present in our benchmark present challenges that can be addressed with a more structured approach to the learning problem, and we hope this stimulates further directions for future research.
V-E Common Failures
(a) highbar
(b) door
(c) hurdle
Figure 8: Failure Scenarios. This figure presents a selection of common failures that occur while training our benchmark tasks.
In this subsection, we remark on notable challenges and common failures for some representative tasks in our benchmark, which denote the challenge in learning with high-dimensional action spaces and limited planning horizon of the state-of-the-art RL algorithms.
Common Failure on highbar. In the highbar task, the Unitree H1 robot conservatively learns to maintain contact with the bar to avoid episode termination, but experiences difficulties in performing the whole-body rotation trajectory. This is indicative of short horizon planning and, despite the availability of dense rewards, is a recurrent challenge in many of the long-horizon benchmark tasks.
Common Failure on door. In the door task, the robot is well-guided to turn the door hatch to unlock the door, but it finds it challenging to learn the precise motion required to pull the door towards its opening position. This is mainly because pulling the door requires not only pulling its arm but also moving the whole body backwards. The coordination between multiple body parts and seamless interaction between manipulation and locomotion skills are common challenges in training humanoid robots.
Common Failure on hurdle. In the hurdle environment, the robot learns to run forward with the expected velocity but does not recognize the need to surpass the hurdle by jumping, which is a hard exploration problem. Previous work has shown that in OpenAI gym Walker2d, the forward-moving reward is sufficient to learn this behavior [25]. On the other hand, the humanoid robot finds conservative poses to collide with the hurdle such that it can stabilize without terminating the episode after hitting the obstacle, without further exploring high-reward jumping behaviors.
VI Conclusion
We presented HumanoidBench, a high-dimensional humanoid robot control benchmark. Ours is the first example of a comprehensive humanoid environment with a diversity of locomotion and manipulation tasks, ranging from toy examples to practical humanoid applications. We set a high bar with our complex tasks, in the hope to stimulate the community to accelerate the development of whole-body algorithms for such robotic platforms.
Future work. HumanoidBench already includes multimodal high-dimensional observations in the form of egocentric vision and whole-body tactile sensing. While our experiments only benchmarked the performance of state-based environments, studying the interplay between different modalities is a compelling direction for future work.
Extensions of the humanoid environment will also eventually include more realistic objects and environments with real-world diversity and higher-quality rendering. As for dexterous manipulation tasks, we envision screwing and furniture assembly tasks being part of our framework, given that they are particularly tailored for bimanual manipulation.
Here we have focused on reinforcement learning algorithms because humanoids are a setting where collecting physical demonstrations may be challenging. However, we believe that other means could be employed to bootstrap learning (e.g., learning from human videos).
Finally, while this was not the focus of our work, the impressive results obtained via domain randomization in the newly developed MuJoCo MJX show promise to study sim-to-real transfer in more depth, following the large success of the field in quadrupedal locomotion [20].
Appendix
[…]
1Table of Contents
Appendix A Additional Components
In addition to the Unitree H1 robot, we also include model files for the Agility Robotics Digit humanoid. We add a custom head, together with egocentric vision and whole-body tactile sensing, similarly to the H1. We provide a torque controlled version of this robot.
We also provide two additional end effectors for the Unitree H1 robot, that is, a Robotiq 2F-45 parallel-jaw gripper, and a dexterous hand available in the Unitree model collection.
Appendix B Simulated Environment Details
B-A Observation Space
The observation space for the robot state (joint positions and velocities) comprises $151$ dimensions, with $49$ dimensions representing the humanoid robot body and $51$ dimensions representing each hand. The observation space can also include environment states for tasks interacting with objects, as described in detail in Section B-E.
B-B Action Space
The action space is the same across all environments. We normalize the action space to be $[-1,1]^{\lVert A\rVert}$, where $\lVert A\rVert=61$ ($19$ for the humanoid body and $21$ for each hand). We use position control for the benchmarking results in this paper.
B-C Simulation Performance
HumanoidBench is based on MuJoCo [52], which provides fast and accurate physics simulation. We provide a diverse set of configurations, such as the full humanoid model with two hands, the humanoid model without hands, and the humanoid model with collision meshes only on its feet. We use the fastest model (i.e. collision meshes only on feet) to train the low-level reaching policies, described in Section C-B.
[…]
Configuration FPS
Without hands $2450$
Simplified body collisions $3600$
Collisions only for feet $5100$
Default $1050$
TABLE II: HumanoidBench Simulation Performance.
B-D Whole-body Tactile Sensing
(a) Refined meshes
(b) Contact before refinement
(c) Contact after refinement
Figure 9: Refinement of collision meshes. As a result of the mesh refinement (see (a), where each colored section indicates a different collision mesh), our model can detect a higher number of contact points, as shown by the yellow disks in the figure. This results in a better spatial discretization of the tactile readings.
We implement whole-body tactile sensing by employing MuJoCo touch grid, which aggregates pressure and shear contact forces into discrete bins. Similar distributed force readings have been captured on real-world systems both on humanoid bodies [36] and end-effectors [46]. To increase the number of contact point candidates and fully exploit the spatial resolution of the touch grid, we subdivide the original meshes into many smaller meshes. Specifically, we build on top of CoACD [54], which in addition makes sure that the resulting meshes are all convex. This procedure results in finer contact resolution when the MuJoCo physics engine computes collisions. An example is depicted in Figure 9, with a larger number of contact points generally resulting in a considerably better discretization of the tactile readings. The full model with refined meshes and tactile readings runs at 550 FPS.
B-E Task Specification
Before enumerating the environment details below, let us define auxiliary functions and variables that are employed in the reward functions of multiple environments:
  • $tol\left(x,(x_{\text{lower}},x_{\text{upper}}),m\right)$ is a function provided in the DeepMind Control Suite package [51]. This is denoted there as a tolerance function that returns $1$ when the evaluated value is within the bounds, i.e., $x\in(x_{\text{lower}},x_{\text{upper}})$, and between 00 and $1$ otherwise. The margin $m$ regulates the slope of the function, i.e., how far from the bounds the function approaches 00.
  • $height\left((x_{\text{lower}},x_{\text{upper}}),m\right)$ is a variable defined as $tol\left(z_{\text{head}},(x_{\text{lower}},x_{\text{upper}}),m\right)$ that rewards the head height, where $z_{\text{head}}$ is the vertical coordinate of the robot head position. Whenever the arguments are not indicated, we assume $x_{\text{lower}}=1.65$, $x_{\text{upper}}=+\infty$, and $m=0.4125$.
  • $d(object_{\text{A}},object_{\text{B}})$ is the distance between object A and object B.
  • $upright\left((x_{\text{lower}},x_{\text{upper}}),m\right)$ is a variable defined as $tol\left(z_{\text{proj}},(x_{\text{lower}},x_{\text{upper}}),m\right)$, which rewards the alignment of the robot torso with respect to the vertical axis, where $z_{\text{proj}}$ is the unit projection of the $z$-axis in the robot body frame onto the $z$-axis in the global frame. Whenever the arguments are not indicated, we assume $x_{\text{lower}}=0.9$, $x_{\text{upper}}=+\infty$, and $m=1.9$.
  • $stand:=height\times upright$ represents a standing posture reward.
  • $e:=0.2\cdot\left[4+\frac{1}{|u|}\sum_{i}tol(u_{i},(0,0),10)\right]$ rewards small control effort, where $u$ is the vector of actuation inputs.
  • $stable=stand\times e$, densely rewards stable standing configurations.
  • $v_{x}$ is the robot velocity in the $x$ direction (positive forward) of the robot body coordinate frame.
  • $v_{y}$ is the robot velocity in the $y$ direction (positive left) of the robot body coordinate frame.
  • $z_{\text{item}}$ is the vertical $z$ coordinate of the indicated ‘item’ in the global frame.
  • $pos_{\text{item}}$ is the 3D position of the indicated ‘item’ in the global frame.
B-E1 walk
[…]
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.2$.
Reward Implementation.
$R(s,a)=stable\times tol(v_{x},(1,+\infty),1).$
B-E2 stand
Objective. Maintain a standing pose.
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. Let:
  • $still_{x}=tol(v_{x},(0,0),2)$
  • $still_{y}=tol(v_{y},(0,0),2)$,
Then:
$R(s,a)=stable\times mean(still_{x},still_{y})$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.2$.
B-E3 run
[…]
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation.
$R(s,a)=stable\times tol(v_{x},(5,\infty),5)$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.2$.
B-E4 sit
Objective. Sit onto a chair situated closely behind the robot.
Observation. In sit_simple, the observation is a vector containing all joint positions on the robot unit. In sit_hard, we allow the chair to be moved, and the robot initial position is randomized, so the observation includes position and orientation of the chair as well.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset. Note that in sit_hard, the robot is rotated at a random angle $\alpha\in[-1.8\ \rm{rad},1.8\ \rm{rad}]$ with position initialized at random values $x\in[0.2,0.4],y\in[-0.15,0.15]$
Reward Implementation. Let
  • $sitting_{x}=tol(x_{robot}-x_{chair},(-0.19,0.19),0.2$
  • $sitting_{y}=tol(y_{robot}-y_{chair},(0,0),0.1$
  • $sitting_{z}=tol(z_{robot},(0.68,0.72),0.2)$
  • $posture=tol(z_{\text{head}}-z_{\text{IMU}},(0.35,0.45),0.3)$
  • $still_{x}=tol(v_{x},(0,0),2)$
  • $still_{y}=tol(v_{y},(0,0),2)$
then, the reward of this task is
$(\displaystyle R(s,a)=($$(0.5\cdot sitting_{z}+0.5\cdot sitting_{x}\times sitting_{y})$
×$\times upright\times posture)$
×$\times e\times mean(still_{x},still_{y})$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.5$.
B-E5 balance
Objective. Balance on the unstable board. There are two variants to this environment: the balance_simple variant’s spherical pivot beneath the board does not move, while the balance_hard variant’s pivot does.
Observation. The observation comprises the joint positions and velocities of the robot and those of the board.
Initialization. The robot is initialized to a standing position on the unstable board, with random noise added to all joint positions during each episode reset.
Reward Implementation. Let
  • $still_{x}=tol(v_{x},(0,0),2)$
  • $still_{y}=tol(v_{y},(0,0),2)$
  • $still=\frac{1}{2}(still_{x}+still_{y})$
  • $height_{\text{robot}}=height((2.15,+\infty))$
Then:
$R(s,a)=(e\times still)\times(height_{\text{robot}}\times upright)$
Termination. The episode terminates after 1000 steps, or when one of the following condition satisfies:
  • $z_{robot}<0.8$
  • The sphere collides with anything other than the floor and the standing board.
  • The standing board collides with the floor.
B-E6 stair
[…]
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. Let
  • $vertical_{\text{foot,left}}=tol(z_{\text{head}}-z_{\text{foot,left}},(1.2,+\infty),0.45)$
  • $vertical_{\text{foot,right}}=tol(z_{\text{head}}-z_{\text{foot,right}},(1.2,+\infty),0.45)$
Then:
$R(s,a)=e$$\times tol(v_{x},(1,+\infty),1)\times upright((0.5,1),1.9)$
×$\times(vertical_{\text{foot,left}}\times vertical_{\text{foot,right}})$
Termination. The episode terminates after 1000 steps, or when $z_{\text{proj}}<0.1$.
B-E7 slide
[…]
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. Let
  • $vertical_{\text{foot,left}}=tol(z_{\text{head}}-z_{\text{foot,left}},(1.2,+\infty),0.45)$
  • $vertical_{\text{foot,right}}=tol(z_{\text{head}}-z_{\text{foot,right}},(1.2,+\infty),0.45)$
Then:
$R(s,a)=e$$\times tol(v_{x},(1,+\infty),1)\times upright((0.5,1),1.9)$
×$\times(vertical_{\text{foot,left}}\times vertical_{\text{foot,right}})$
Termination. The episode terminates after 1000 steps, or when $z_{\text{proj}}<0.6$.
B-E8 pole
Objective. Travel in forward direction over a dense forest of high thin poles, without colliding with them.
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. Let
  • $\gamma_{\text{collision}}=\begin{cases}0.1,&,\text{robot collides with pole}\\ 1,&\text{otherwise}\end{cases}$
Then,
$R(s,a)=\gamma_{\text{collision}}\times(0.5\cdot stable+0.5\cdot tol(v_{x},(1,+\infty),1)).$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.6$.
B-E9 reach
Objective. Reach a randomly initialized 3D point with the left hand.
Observation. Joint positions and velocities of the robot, left hand position of robot, and reaching target position.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. Let
  • $d_{\text{hand}}=d(hand_{\text{left}},goal)$
  • $health=5\cdot z_{pelvis,proj}$
  • $penalty_{motion}=\|v_{robot}\|^{2}$
  • $close=\begin{cases}5,&d_{\text{hand}}<1\\ 0,&\text{otherwise}\end{cases}$
  • $success=\begin{cases}10,&d_{\text{hand}}<0.05\\ 0,&\text{otherwise}\end{cases}$
Then,
$R(s,a)=-10^{-4}\cdot penalty_{motion}+health+close+success$
Termination. The episode terminates after 1000 steps.
B-E10 hurdle
[…]
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. Let
$\gamma_{\text{collision}}=\begin{cases}0.1&,\text{robot is colliding with wall}\\ 1&,\text{otherwise}\end{cases}$
Then, we formulate the reward of this task as:
$R(s,a)=stable\times tol(v_{x},(5,\infty),5)\times\gamma_{\text{collision}}$
Termination. The episode terminates after 1000 steps.
B-E11 crawl
[…]
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. Let:
  • $height_{\text{crawl}}=height((0.6,1),1)$
  • $height_{\text{IMU}}=tol(z_{\text{IMU}},(0.6,1),1)$
  • $quat_{\text{crawl}}=\begin{bmatrix}0.75&0&0.65&0\end{bmatrix}$ is the expected quaternion expected when the robot is crawling.
  • $orientation=tol(\|quat_{\text{pelvis}}-quat_{\text{crawl}}\|,(0,0),1)$ rewards correct robot body orientation.
  • $tunnel=tol(y_{\text{IMU}},(-1,1),0)$ rewards the robot when its $y$ coordinate is inside the tunnel.
  • $speed=tol(v_{x},(1,+\infty),1)$
Then, the reward is formulated as:
$R=tunnel\times(0.1\cdot e$ $+0.25\cdot\min(height_{\text{crawl}},height_{\text{IMU}})$
$+0.25\cdot orientation+0.4\cdot speed)$
Termination. The episode terminates after 1000 steps.
B-E12 maze
Objective. Reach the goal position in a maze by taking multiple turns at the intersections.
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. For each checkpoint in the maze, we assign $v_{target}$ to be the velocity prescribed to travel from the previous checkpoint towards next one. Then, let
  • $move=$
    $tol(v_{x}-v_{target_{x}},(0,0),v_{target_{x}})$
    ×$\times tol(v_{y}-v_{target_{y}},(0,0),v_{target_{y}})$
  • $\gamma_{\text{collision}}=\begin{cases}0.1&,\text{robot is colliding with wall}\\ 1&,\text{otherwise}\end{cases}$
  • $proximity=tol(d(checkpoint,pos_{robot}),(0,0),1)$
We formulate the reward as
$R(s,a)=($ $0.2\cdot stable+0.4\cdot move$
$+0.4\cdot proximity)\times\gamma_{\text{collision}}$
We also provide a sparse reward of $i\times 100$ for arriving at the $i^{th}$ checkpoint.
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.2$.
B-E13 push
Objective. Move a box to a randomly initialized 3D point on a table.
Observation. Joint positions and velocities of the robot, left hand position of the robot, box destination, box position, and box velocity.
Initialization. The robot is initialized to a standing position. The box and its destination are initialized at a random location on the table. Random noise is added to all joint positions during each episode reset.
Reward Implementation. Let:
  • $d_{\text{goal}}=d(box,destination)$
  • $success=\mathbbm{1}_{d_{\text{goal}}<0.05}$
  • $d_{\text{hand}}=d(box,hand_{\text{left}})$
Then the reward is
$R(s,a)=\alpha_{s}\cdot success-\alpha_{t}\cdot d_{\text{goal}}-\alpha_{h}\cdot d_{\text{hand}}$
where by default $\alpha_{s}=1000,\alpha_{t}=1,\alpha_{h}=0.1$.
Termination. The episode terminates after 500 steps, or when $d_{\text{goal}}<0.05$.
B-E14 cabinets
Objective. Open four different types of cabinet doors (e.g. hinge door, sliding door, drawer) and perform different manipulations (see subtasks below) for objects inside the cabinet.
Observation. Positions and velocities of the robot’s joints, the cabinets’ joints, and the objects situated inside the cabinets.
Initialization. The robot is initialized to a standing position, with random noise added to all joint positions during each episode reset.
Reward Implementation. The reward of this task changes based on the occurring subtask.
Subtask 1 is to open the sliding door (second highest one), with reward
$R_{1}=0.2\cdot stable+0.8\cdot|l_{\text{cabinet}}/0.4|$
where the cabinet joint position $l_{\text{cabinet}}$ is a value in $[0,0.4]$.
Subtask 2 is to open the drawer (the lowest one), with reward
$R_{2}=0.2\cdot stable+0.8\cdot|l_{\text{drawer}}/0.45|$
where the drawer joint position $l_{\text{drawer}}$ is a value in $[0,0.45]$.
Subtask 3 is to put the cube from the drawer into the hinge-based cabinet (the third highest one). Both the left and right hinge-based cabinet doors’ joint positions $\alpha_{\text{door,left}},\alpha_{\text{door,right}}$ are values in $[0,1.57]$. Let
  • $open_{\text{door,left}}=\min(1,|\alpha_{\text{door,left}}|)$
  • $open_{\text{door,right}}=\min(1,|\alpha_{\text{door,right}}|)$
  • $d_{\text{destination,x}}=tol(x_{cube}-0.9,(-0.3,0.3),0.3)$
  • $d_{\text{destination,y}}=tol(y_{cube},(-0.6,0.6),0.3)$
  • $d_{\text{destination,z}}=tol(z_{cube}-0.94,(-0.15,0.15),0.3)$
  • $r_{\text{destination}}=0.3\cdot mean(d_{\text{destination,x}},d_{\text{destination,y}})+0.7\cdot d_{\text{destination,z}}$
Then the reward of this subtask, $R_{3}$, is formulated as follows
$r_{3}$ $=0.5\cdot\max(open_{\text{door,left}},open_{\text{door,right}})+0.5\cdot r_{\text{destination}}$
$R_{3}$ $=0.2\cdot stable+0.8\cdot r_{3}$
Subtask 4 is to put the original cube from the hinge-based left-right cabinet (third highest) into the pull-up cabinet (highest). The pull-up door has a joint position $\alpha_{\text{pull}}$ in $[0,1.57]$. Let
  • $open_{\text{pull}}=\min(1,|\alpha_{\text{pull}}|)$
  • $d_{\text{destination,x}}=tol(x_{cube}-0.9,(-0.3,0.3),0.3)$
  • $d_{\text{destination,y}}=tol(y_{cube},(-0.6,0.6),0.3)$
  • $d_{\text{destination,z}}=tol(z_{cube}-1.54,(-0.15,0.15),0.3)$
  • $r_{\text{destination}}=0.3\cdot mean(d_{\text{destination,x}},d_{\text{destination,y}})+0.7\cdot d_{\text{destination,z}}$
Then the reward of this subtask, $R_{4}$, is formulated as follows
$r_{4}=$ $0.5\cdot open_{\text{pull}}+0.5\cdot r_{\text{destination}}$
$R_{4}=$ $0.2\cdot stable+0.8\cdot r_{4}$
The reward during the occurrence of subtask $i$ is $R_{i}$. Upon the completion of subtask $i$, a sparse reward of $i*100$ is offered. At the timestep where all subtasks are completed, a sparse reward of $1000$ is added to the total reward.
Termination. The episode terminates after 1000 steps or whenever all subtasks are complete.
B-E15 highbar
Objective. Athletically swing while staying attached to a horizontal high bar until reaching a vertical upside-down position.
Observation. Joint positions and velocities of the robot.
Initialization. The robot is initialized such that its hand is gripping the high bar, and its body is hanging from it.
Reward Implementation. Let
  • $upright_{highbar}=upright((-\infty,-0.9),1.9)$
  • $feet=tol((z_{\text{foot,left}}+z_{\text{foot,right}})/2,(4.8,+\infty),2)$
Then,
$R(s,a)=upright_{highbar}\times feet\times e$
Termination. The episode terminates after 1000 steps, or when $z_{\text{head}}<2$.
B-E16 door
Objective. Pull a door open using its doorknob, and traverse through the doorpath while keeping the door open.
Observation. Joint positions and velocities of the robot, door hinge, and door hatch.
Initialization. The robot is initialized to a standing position. The door hinge joint and door latch are initialized such that the door is completely closed. For the current implementation, the door can only be pulled towards the robot, limiting its hinge joint position to be within $[0,1.4]$. The door latch has a joint range of $[0,2]$ (being pulled downwards until $2$ radians from positive x-axis, clockwise). Random noise is added to all joint positions during each episode reset.
Reward Implementation. Let
  • $open_{\text{door}}=\min(1,q_{\text{door}}^{2})$
  • $open_{\text{hatch}}=tol(q_{\text{hatch}},(0.75,2),0.75)$
  • $proximity_{\text{door}}=$
    $tol($ $\min(d(hand_{\text{left}},door),d(hand_{\text{right}},door)),$
    (0,0.25),1(0,0.25),1)( 0 , 0.25 ) , 1 )
  • $passage=tol(x_{\text{IMU}},(1.2,+\infty),1)$
Then, the reward of this task is:
$R$ $=0.1\cdot stable+0.45\cdot open_{\text{door}}+0.05\cdot open_{\text{hatch}}$
$+0.05\cdot proximity_{\text{door}}+0.35\cdot passage$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.58$.
B-E17 truck
Objective. Unload packages from a truck by moving them onto a platform.
Observation. Joint positions and velocities of the robot, and positions and velocities of packages.
Initialization. The robot is initialized to a standing position. Packages are initialized to be on the truck. Random noise is added to all joint positions during each episode reset.
Reward Implementation. The reward of this task relies on the subsets of packages based on three categories: (1) being on truck ($p_{\text{truck}}$), (2) being picked up ($p_{\text{picked}}$), and (3) being on table ($p_{\text{table}}$). Let
  • $truck=tol(\min_{p\in p_{\text{truck}}}\|pos_{p}-pos_{\text{pelvis}}\|,(0,0.2),4)$
  • $picked=tol(\min_{p\in p_{\text{picked}}}\|pos_{p}-pos_{\text{pelvis}}\|,(0,0.2),4)$
  • $table=tol(\min_{p\in p_{\text{table}}}\|pos_{p}-pos_{\text{table}}\|,(0,0.2),4)$
  • $r_{location}=100\cdot(p_{table}+p_{picked}-p_{truck})$
Then, the reward is:
$R(s,a)=r_{location}+upright\times(1+truck+picked+table)$
If all packages are picked up, an additional sparse reward of $1000$ is provided, and the episode terminates thereafter.
Termination. The episode terminates after 1000 steps, or when all packages are delivered onto the table.
B-E18 cube
Objective. Manipulate two cubes, each cube in one hand, until they both correspond with a specific, randomly initialized target orientation.
Observation. Joint positions and velocities of the robot and the two cubes to be manipulated on hand. The transparent cube in front of the robot is an indication of the target orientation for in-hand cubes, and its orientation is also in the state.
Initialization. The robot is initialized to a standing position. The cubes are initialized at a random orientation right above the robot hands. Random noise is added to all joint positions during each episode reset.
Reward Implementation. Let
  • $still_{x}=tol(v_{x},(0,0),2)$
  • $still_{y}=tol(v_{y},(0,0),2)$
  • $still=mean(still_{x},still_{y})$
  • $quat_{\text{target}}$ denotes the target orientation for both in-hand cubes.
  • $orientation=$
    $\frac{1}{2}[(quat_{\text{cube,left}}-quat_{\text{target}})^{2}+(quat_{\text{cube,right}}-quat_{\text{target}})^{2}]$
  • $proximity_{\text{cube}}=$
    $\frac{1}{2}[$ $tol(d(cube_{\text{left}},hand_{\text{left}}),(0,0),0.5)$
    $+tol(d(cube_{\text{right}},hand_{\text{right}}),(0,0),0.5)]$
Then the reward of this task is:
$R=$ $0.2\cdot(stable\times still)$
$+0.5\cdot orientation+0.3\cdot proximity_{\text{cube}}$
Termination. The episode terminates after 500 steps, or when $z_{\text{pelvis}}<0.5$, $z_{\text{cube,left}}<0.5$, or $z_{\text{cube,right}}<0.5$.
B-E19 bookshelf
Objective. The bookshelf environment mainly concerns relocating five objects across the various shelves. There are five designated subtasks, resembling five different relocations (each for a different item and destination location). The items involved in the subtasks are colored from brightest to darkest in a red shade, where the brighter shade of red shows that the object’s relocation is an earlier subtask to complete. The subtasks must be completed in order. The order is always the same in bookshelf_simple, while it is randomized at every episode in bookshelf_hard.
Observation. Positions and velocities of the robot’s joints and all objects on the bookshelf, including those that are not associated with any subtasks, as well as an index representing the next object to be relocated.
Initialization. The robot is initialized to a standing position. All objects on the bookshelf are currently initialized at a default position. Random noise is added to all joint positions during each episode reset. In the simple version of this environment, the objects to move on the bookshelf and their destinations are fixed; in the hard version, both of them are randomized at the beginning of every episode instead.
Reward Implementation. For each subtask $t_{i}$ where $i\in\{1,2,\dots,5\}$, its reward $r_{i}$ is formulated as follows. First, let
  • $proximity_{\text{destination}}=$
    $tol(d(object,destination),(0,0.15),1)$
  • $d_{\text{hand}}=$
    $\min(d(object,hand_{\text{left}}),d(object,hand_{\text{right}}))$
  • $proximity_{\text{hand}}=\exp(-d_{\text{hand}})$.
Then, the subtask reward is
$r_{i}=$ $0.4\cdot proximity_{\text{hand}}$
$+0.2\cdot stable+0.4\cdot proximity_{\text{destination}}$
and the reward for that step is the subtask reward per se. The current subtask is considered complete when the distance between the destination and object of current subtask is less than $0.15$.
An additional sparse reward of $100*i$ is added to the timestep of subtask $i$’s completion.
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.58$, $z_{\text{object}}<0.5$, or all subtasks succceeded.
B-E20 basketball
Objective. Catch a ball coming from random direction and throw it into the basket.
Observation. Positions and velocities of the robot’s joints and the basketball.
[…]
Reward Implementation. The task is divided into two stages:
$\begin{cases}\text{catch}:&\text{before the basketball collides with anything}\\ \text{throw}:&\text{after the basketball experiences one collision}\end{cases}$
The rewards at different stages are formulated differently. Let
  • $proximity_{\text{hand}}=$
    $tol($ $\max(d(ball,hand_{\text{left}}),d(ball,hand_{\text{right}})),$
    (0,0.2),1(0,0.2),1)( 0 , 0.2 ) , 1 )
  • $aim=tol(d(ball,basket),(0,0),7)$.
Then, the reward at each stage of the task is formulated as follows:
$R_{\text{catch}}(x,u)$ $=0.5\cdot proximity_{\text{hand}}+0.5\cdot stable$
$R_{\text{throw}}(x,u)$ $=0.05\cdot proximity_{\text{hand}}$
$+0.15\cdot stable+0.8\cdot aim$
At the earliest timestep where $d(ball,basket)\leq 0.05$, the episode terminates, and a large sparse reward of $1000$ is given.
Termination. The episode terminates after 500 steps, or when $z_{\text{pelvis}}<0.5$, $z_{\text{ball}}<0.5$, or when success has been achieved.
B-E21 window
Objective. Grab a window wiping tool and keep its tip parallel to a window by following a prescribed vertical speed (in absolute value).
Observation. Joint positions and velocities of the robot and the window wiping tool (position and velocities of the entire tool, in addition to those of the rotational joint attached between the wipe and the rod of the tool).
Initialization. The robot is initialized to a standing position. The window wiping tool is initialized above the robot unit’s hand in a parallel direction to the window frames. Random noise is added to all joint positions during each episode reset.
Reward Implementation. Let
  • $proximity_{\text{tool}}=$
    $\frac{1}{2}[tol(d(tool,hand_{\text{left}}),(0,0),0.5)$
    $+tol(d(tool,hand_{\text{right}}),(0,0),0.5)]$
  • $d_{\text{window}}=tol(d(head,window),(0.4,0.4),0.1)$
  • $move_{\text{wipe}}=tol(|v_{\text{wipe,z}}|,(0.5,0.5),0.5)$
The manipulation reward is defined as:
$r_{\text{manipulation}}=$ $0.4\cdot move_{\text{wipe}}+0.4\cdot proximity_{\text{tool}}$
$+0.2\cdot(stable\times d_{\text{window}})$
Meanwhile, five sites are put on the four corners and center of the wipe to detect coverage of contact. Let these sites be $s_{1},s_{2},\dots,s_{5}$, then the reward for contact between the tool and window is defined as:
$r_{\text{contact}}$ $=\frac{1}{5}\sum_{s_{i}}tol(x_{s_{i}},(0.92,0.92),0.4)$
Then, the reward of this task is:
$R=0.5\cdot r_{\rm manipulation}+0.5\cdot r_{\rm contact}$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.58$ or $z_{\text{tool}}<0.58$.
B-E22 spoon
Objective. Grab a spoon and use it to follow a circular pattern inside a pot.
Observation. Joint positions and velocities of the robot and the position and velocity of the spoon, as well as the target position that the spoon should be at the current timestep.
Initialization. The robot is initialized to a standing position. The spoon is initialized at a specified position on the table, leftwards of the pot. Random noise is added to all joint positions during each episode reset.
Reward Implementation. Let
  • $t:=$ current timestep
  • $proximity_{\text{tool}}=$
    $\frac{1}{2}[tol(d(tool,hand_{\text{left}}),(0,0),0.5)$
    $+tol((tool,hand_{\text{right}}),(0,0),0.5)]$
  • $destination=\begin{bmatrix}x_{\text{pot}}+0.06\cos(\frac{t\pi}{20})\\ y_{\text{pot}}+0.06\sin(\frac{t\pi}{20})\\ z_{\text{pot}}\end{bmatrix}$
  • $r_{\text{trajectory}}=tol(d(spoon,destination),(0,0),0.15)$
  • $r_{\text{destination}}=$
    $\frac{1}{3}\sum_{i\in\{x,y,z\}}\mathbbm{1}_{i_{\text{spoon}}\text{in the pot}}$
Then:
$R(s,a)$ $=0.15\cdot stable+0.25\cdot proximity_{tool}$
$+0.25\cdot r_{\text{destination}}+0.35\cdot r_{\text{trajectory}}$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.58$.
B-E23 kitchen
Objective. Execute a sequence of actions in a kitchen environment, namely, open a microwave door, move a kettle, and turn burner and light switches.
Observation. Joint positions and velocities of the robot, and positions and velocities of kitchenware.
Initialization. The robot is initialized to a standing position. Kitchenware is initialized at specified positions. Random noise is added to all joint positions during each episode reset.
Reward Implementation. A subtask in kitchen is considered complete if the distance between an object and its goal position is lower than a specified threshold. The sparse reward is the number of subtasks completed.
Termination. The episode terminates after 500 steps.
B-E24 package
Objective. Move a box to a randomly initialized target position, also tested in the ablation in Figure 7.
Observation. Joint positions and velocities of the robot, both hand positions of the robot, package destination, package position, and package velocity.
Initialization. The robot is initialized to a standing position. The package position and its destination are randomly initialized within a specific area. Random noise is added to all joint positions during each episode reset.
Reward Implementation. Let
  • $height_{\text{package}}=\min(1,z_{\text{package}})$
  • $success=\mathbbm{1}_{d(package,destination)<0.1}$
  • $d_{\text{hand}}=d(package,hand_{\text{left}})+d(package,hand_{\text{right}})$.
Then, the reward is:
$R(s,a)=$ $-3\cdot d(package,destination)-0.1\cdot d_{\text{hand}}$
$+stable+height_{\text{package}}+1000\cdot success$
Termination. The episode terminates after 1000 steps, or when $d(package,destination)<0.1$.
B-E25 powerlift
Objective. Lift a barbell of a designated mass.
Observation. Joint positions and velocities of the robot and barbell.
Initialization. The robot is initialized to a standing position. The barbell is initialized on the ground. Random noise is added to all joint positions during each episode reset.
Reward Implementation. Let
  • $height_{barbell}=tol(z_{barbell},(1.9,2.1),2)$.
Then,
$R(s,a)=0.2\cdot stable+0.8\cdot height_{barbell}.$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.2$.
B-E26 room
[…]
Observation. Joint positions and velocities of the robot and scattered objects.
Initialization. The robot is initialized to a standing position. The scattered objects’ positions are randomly initialized within a specific area. Random noise is added to all joint positions during each episode reset.
Reward Implementation. Let
  • $X$ be a matrix containing the location of all scattered objects in 3D coordinates.
  • $cleanness=tol(\max(Var(X_{:,0}),Var(X_{:,1})),(0,0),3)$, where $Var(X_{:,0})$ is the variance of x-coordinates of all objects’ locations, and $Var(X_{:,1})$ is such variance for y-coordinates.
Then,
$R(s,a)=0.2\cdot stable+0.8\cdot cleanness$
Termination. The episode terminates after 1000 steps, or when $z_{\text{pelvis}}<0.3$.
B-E27 insert
Objective. Insert the ends of a rectangular block into two small pegs. Two versions, insert_small and insert_normal present different object sizes.
Observation. Joint positions and velocities of the robot, rectangular block to insert, and two provided small pegs.
Initialization. The robot is initialized to a standing position. The rectangular block and two pegs are initialized on the table at specific orientation and position.
Reward Implementation. Let the rectangular blocks be formualted to have two ends $end_{a}$, $end_{b}$, which must be respectively inserted to the pegs $peg_{a}$, $peg_{b}$.
  • $proximity_{peg,site}=tol(d(peg,site),(0,0),0.5)$
  • $proximity_{block}=$
    $mean(proximity(peg_{a},end_{a}),proximity(peg_{b},end_{b}))$
  • $height(peg)=tol(z_{peg}-1.1,(0,0),0.15)$
  • $height_{pegs}=mean(height(peg_{a}),height(peg_{b}))$
  • $proximity_{hands}=$
    $mean($ $proximity(peg_{a},hand_{left}),$
    $proximity(peg_{b},hand_{right}))$
The reward of this task is phrased as:
$R(s,a)=$ $(0.5\cdot stable+0.5\cdot proximity_{block})$
×$\times(0.5\cdot height_{pegs}+0.5\cdot proximity_{hands})$
Termination. The episode terminates after 1000 steps, or when any of the blocks or pegs are at a height lower than $0.5$ from floor.
Appendix C Training Details
C-A Baseline Implementation Details
For SAC, we use the implementation from JaxRL Minimal [13]. For PPO, we use the Stable-Baselines3 [44] implementation with 4 parallel environments. For DreamerV3 and TD-MPC2, we use their official code. For DreamerV3, we use the ‘medium’ configuration, with an update-to-data ratio of 64. For TD-MPC2, we use the 5M configuration. Unless specified, we use their default hyperparameters.
C-B Reaching Policy Implementation Details
We train the low-level reaching policies described in Section V-D using PPO, largely parallelized on GPU using MuJoCo MJX. We employ the PureJaxRL [30] implementation, using their default parameters, except a higher number of environments as detailed in Section V-D, 16 steps per environment, and an entropy coefficient of 0.001.
C-C Benchmarking Results
[…]
DreamerV3 TD-MPC2 SAC Target
walk $800.2\pm 158.7$ $782.0\pm 109.2$ $31.7\pm 24.0$ $700.0$
stand $622.7\pm 404.8$ $809.0\pm 137.1$ $208.3\pm 105.6$ $800.0$
run $633.8\pm 222.4$ $93.3\pm 14.3$ $5.0\pm 2.1$ $700.0$
reach $7580.9\pm 1951.0$ $7316.1\pm 2112.1$ $4565.1\pm 212.8$ $12000.0$
hurdle $126.2\pm 59.4$ $46.4\pm 10.8$ $13.2\pm 8.8$ $700.0$
crawl $878.8\pm 122.7$ $957.4\pm 17.5$ $330.0\pm 111.9$ $700.0$
maze $272.3\pm 116.6$ $244.3\pm 97.7$ $144.8\pm 17.8$ $1200.0$
sit_simple $891.4\pm 38.4$ $411.1\pm 368.0$ $148.3\pm 103.8$ $750.0$
sit_hard $433.4\pm 355.9$ $343.0\pm 381.7$ $55.0\pm 18.2$ $750.0$
balance_simple $19.8\pm 7.0$ $40.5\pm 23.9$ $61.5\pm 1.1$ $800.0$
balance_hard $45.9\pm 27.4$ $48.2\pm 28.5$ $42.5\pm 22.6$ $800.0$
stair $131.1\pm 43.6$ $70.4\pm 7.1$ $14.1\pm 6.8$ $700.0$
slide $436.5\pm 200.1$ $119.0\pm 35.9$ $6.3\pm 2.8$ $700.0$
pole $658.3\pm 343.3$ $226.3\pm 116.1$ $46.3\pm 26.4$ $700.0$
push $-1251.9\pm 659.8$ $-258.7\pm 66.5$ $-97.9\pm 147.0$ $700.0$
cabinet $57.3\pm 66.3$ $112.8\pm 142.9$ $211.8\pm 33.8$ $2500.0$
highbar $8.9\pm 5.8$ $0.3\pm 0.0$ $9.4\pm 3.7$ $750.0$
door $213.0\pm 149.3$ $274.7\pm 12.5$ $39.4\pm 25.2$ $600.0$
truck $1103.8\pm 232.9$ $1132.6\pm 72.1$ $1077.5\pm 95.0$ $3000.0$
cube $111.2\pm 59.9$ $54.7\pm 33.1$ $130.7\pm 30.5$ $370.0$
bookshelf_simple $840.4\pm 5.6$ $136.2\pm 71.6$ $346.9\pm 231.5$ $2000.0$
bookshelf_hard $530.2\pm 302.5$ $37.0\pm 1.3$ $293.9\pm 121.6$ $2000.0$
basketball $19.3\pm 2.5$ $42.0\pm 14.8$ $22.1\pm 3.2$ $1200.0$
window $461.0\pm 252.8$ $87.1\pm 37.5$ $62.9\pm 83.8$ $650.0$
spoon $349.7\pm 46.2$ $77.9\pm 80.6$ $87.7\pm 80.5$ $650.0$
kitchen $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $4.0$
package $-18015.2\pm 9477.7$ $-3655.6\pm 1055.0$ $-6718.3\pm 607.0$ $1500.0$
powerlift $315.9\pm 16.9$ $99.1\pm 47.3$ $81.8\pm 46.7$ $800.0$
room $120.5\pm 71.4$ $131.4\pm 56.7$ $12.0\pm 4.9$ $400.0$
insert_small $184.8\pm 26.3$ $129.8\pm 51.9$ $10.8\pm 13.4$ $350.0$
insert_normal $171.5\pm 33.2$ $237.6\pm 9.1$ $46.3\pm 63.1$ $350.0$
TABLE III: Average returns for HumanoidBench. Each number represents average return@10M (return@2M) with the standard deviation for DreamerV3 and SAC (TD-MPC2).
DreamerV3 TD-MPC2 SAC Target
walk $932.4\pm 0.3$ $900.3\pm 47.6$ $68.7\pm 27.0$ $700.0$
stand $932.9\pm 1.1$ $925.7\pm 2.5$ $809.9\pm 194.5$ $800.0$
run $895.9\pm 6.0$ $226.7\pm 23.3$ $104.8\pm 7.4$ $700.0$
reach $9831.6\pm 115.9$ $9727.6\pm 48.9$ $7169.9\pm 874.1$ $12000.0$
hurdle $396.8\pm 39.7$ $196.7\pm 30.5$ $78.6\pm 32.8$ $700.0$
crawl $985.3\pm 0.6$ $985.2\pm 0.4$ $626.5\pm 29.1$ $700.0$
maze $592.5\pm 49.0$ $444.9\pm 22.1$ $269.9\pm 37.8$ $1200.0$
sit_simple $935.7\pm 5.5$ $928.4\pm 1.5$ $842.7\pm 50.8$ $750.0$
sit_hard $914.6\pm 1.5$ $906.3\pm 6.0$ $214.0\pm 47.9$ $750.0$
balance_simple $95.4\pm 8.3$ $95.3\pm 3.1$ $80.7\pm 3.7$ $800.0$
balance_hard $114.0\pm 12.3$ $122.2\pm 13.1$ $71.0\pm 7.9$ $800.0$
stair $411.4\pm 9.7$ $251.9\pm 9.1$ $42.8\pm 0.7$ $700.0$
slide $928.4\pm 2.0$ $311.9\pm 15.1$ $41.4\pm 2.4$ $700.0$
pole $952.2\pm 10.3$ $644.9\pm 21.2$ $440.0\pm 88.4$ $700.0$
push $1000.0\pm 0.0$ $1000.0\pm 0.0$ $352.8\pm 31.5$ $700.0$
cabinet $722.6\pm 7.3$ $721.6\pm 25.9$ $485.9\pm 137.2$ $2500.0$
highbar $83.1\pm 4.6$ $0.9\pm 0.4$ $40.8\pm 41.7$ $750.0$
door $335.7\pm 14.8$ $310.6\pm 10.6$ $251.2\pm 9.0$ $600.0$
truck $1674.3\pm 52.6$ $1457.2\pm 24.3$ $1387.5\pm 10.0$ $3000.0$
cube $237.9\pm 3.4$ $241.1\pm 1.2$ $203.5\pm 27.2$ $370.0$
bookshelf_simple $849.6\pm 0.3$ $825.0\pm 9.6$ $766.5\pm 10.3$ $2000.0$
bookshelf_hard $867.8\pm 8.4$ $320.4\pm 58.9$ $681.5\pm 10.3$ $2000.0$
basketball $808.8\pm 340.5$ $1055.3\pm 4.1$ $192.3\pm 45.6$ $1200.0$
window $765.6\pm 38.4$ $201.1\pm 91.5$ $128.6\pm 170.8$ $650.0$
spoon $421.5\pm 2.5$ $403.5\pm 3.2$ $297.5\pm 26.5$ $650.0$
kitchen $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $4.0$
package $1009.2\pm 4.1$ $1003.3\pm 3.4$ $-3552.8\pm 361.3$ $1500.0$
powerlift $338.6\pm 0.1$ $264.7\pm 23.9$ $171.2\pm 3.6$ $800.0$
room $420.8\pm 51.1$ $353.3\pm 41.5$ $52.2\pm 3.9$ $400.0$
insert_small $239.8\pm 4.6$ $226.4\pm 10.8$ $72.7\pm 21.9$ $350.0$
insert_normal $279.9\pm 9.9$ $273.0\pm 5.7$ $135.3\pm 51.5$ $350.0$
TABLE IV: Maximum returns for HumanoidBench. Each number represents maximum return@10M (return@2M) with the standard deviation for DreamerV3 and SAC (TD-MPC2).